Space vehicles traveling through Earth's radiation belts can intercept a significant number of electrons. The resulting charge buildup can damage electronic components and disrupt operations. Suppose a spherical metallic satellite 1.3 m in diameter accumulates -2.7
Answer #1

a)Surface charge density=Q/A=-2.7*10-6/4*π*0.652=-5.09*10-7C/m2=-0.509μC/m2

b)E=kq/r2=9*109*-2.7*10-6/0.652=-57515N/C